ABC,. Inc just paid a dividend of $12.46. The dividends are expected to grow by 20% in Year 1, 14% in Year 2, and 9% in Year 3. After that, the dividends are expected to grow by 6% each year. If the required rate of return is 12%, what is today's price of the stock?

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

Year 1 dividend = 12.46 (1 + 20%) = 14.952

Year 2 dividend = 14.952 (1 + 14%) = 17.04528

Year 3 dividend = 17.04528 (1 + 9%) = 18.579355

Year 4 dividend = 18.579355 (1 + 6%) = 19.694117

Value at year 3 = D4 / required rate - growth rate

Value at year 3 = 19.694117 / 0.12 - 0.06

Value at year 3 = 19.694117 / 0.06

Value at year 3 = 328.23528

Price of stock = 14.952 / (1 + 0.12)1 + 17.04528 / (1 + 0.12)2 + 18.579355 / (1 + 0.12)3 + 328.23528 / (1 + 0.12)3

Price of stock = $273.79